OpenStudy (anonymous):

show steps please Solve the system by using elimination or elimination with multiplication. x - y = 3 and x + y = 7 Solve the system by using elimination or elimination with multiplication. 2x + 3y = 12 and 3x + 2y = 13

OpenStudy (anonymous):

x-y=3 x+y=7 Add them(elimination) 2x=10 x=5 x-y=3 y=2 (5,2)

OpenStudy (anonymous):

first u need to find the value of either x or y example: x - y = 3 and x + y = 7 x - y=3 x =3+y or y = x-3

OpenStudy (anonymous):

is (5,2) he anwser for the first one?

OpenStudy (anonymous):

yep!

OpenStudy (anonymous):

try to do the second one urself

OpenStudy (anonymous):

2x+3y=12 3x+2y=13 multiply the first one by -2. multiply the second one by 3. -4x-6y=-24 9x+6y=39 5x=15 x=3 Use this hint above if you need it :)

OpenStudy (anonymous):

3, 2?

OpenStudy (anonymous):

good!

OpenStudy (anonymous):

Solve the system by using elimination or elimination with multiplication. x + y = 16 and x + y = -2 is this one no solution?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

3x - 2y = 9 and -3x + 2y = -9 and is this on no solution or infinte solutions??

OpenStudy (anonymous):

notice if you eliminate it, there's nothing left

OpenStudy (anonymous):

k thanks so much. thats what i thought. ya just making sure.

OpenStudy (anonymous):

the last one you asked is infinite

OpenStudy (anonymous):

they're the same thing

OpenStudy (anonymous):

ok good thats what i thought! not on hw they arent lol

OpenStudy (anonymous):

haha true :P

OpenStudy (anonymous):

\[6\sqrt{2}-3\sqrt{2}\]

OpenStudy (anonymous):

3 sqrt 2

OpenStudy (anonymous):

haha i already got it :p this one is hard. one sec.

OpenStudy (anonymous):

\[5\sqrt{3}+8\sqrt{3}-6\sqrt{2}+4\sqrt{2}\]

OpenStudy (anonymous):

13 sqrt3 -2 sqrt 2

OpenStudy (anonymous):

lol, is this school hw?

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

how did you get that?

OpenStudy (anonymous):

\[\sqrt{2 } =x \] \[\sqrt{3} = y\] 5y+8y -6x+4x= 13y -2x and u will get : \[13\sqrt{3}-2\sqrt{2}\]

OpenStudy (anonymous):

that is the easier way to understand it